# Optimization Modeling via Semantic Anchored Alignment

### v1

@@ Abstract
Large language models (LLMs) have opened new paradigms in optimization modeling by enabling the generation of executable solver code from natural language descriptions. Despite this promise, existing approaches typically remain solver-driven: they rely on single-pass forward generation and apply limited post-hoc fixes based on solver error messages, leaving undetected semantic errors that silently produce syntactically correct but logically flawed models. To address this challenge, we propose SAC-Opt, a backward-guided correction framework that grounds optimization modeling in problem semantics rather than solver feedback. At each step, SAC-Opt aligns the original semantic anchors with those reconstructed from the generated code and selectively corrects only the mismatched components, driving convergence toward a semantically faithful model. This anchor-driven correction enables fine-grained refinement of constraint and objective logic, enhancing both fidelity and robustness without requiring additional training or supervision. Empirical results on seven public datasets demonstrate that SAC-Opt improves average modeling accuracy by 7.8%, with gains of up to 21.9% on the ComplexLP dataset. These findings highlight the importance of semantic-anchored correction in LLM-based optimization workflows to ensure faithful translation from problem intent to solver-executable code.
@@ 1 Introduction
Optimization problems arise across domains such as logistics, healthcare, and finance, supporting tasks like planning, allocation, and portfolio optimization (Antoniou & Lu, 2007; Singh, 2012). These problems are typically formulated as mathematical programs and solved using external solvers such as Gurobi (Bixby, 2007), CPLEX (Cplex, 2009), Pyomo (Hart et al., 2011), or COPT Ge et al. (2023). However, translating real-world scenarios into solver-executable code often requires collaboration between domain experts and engineers. This process is time-consuming, hard to scale, and largely inaccessible to non-experts, as reflected by a survey showing that 81% of Gurobi users hold advanced degrees, with nearly half specializing in operations research (Optimization, 2023).
@@
(1) We introduce SAC-Opt, the first optimization modeling framework that establishes a semantics-driven correction method, moving beyond solver-driven syntactic checks. (2) We propose a backward, semantic anchor-guided correction mechanism that progressively aligns models with problem intent, achieving convergence through fine-grained refinement rather than trial-and-error regeneration. (3) We evaluate SAC-Opt on seven public datasets and show that it improves modeling accuracy by 7.8% on average, with a 21.9% gain on the challenging ComplexLP dataset.
@@ 3.5 Iterative Semantic Alignment and Correction
LLM-based Verification: (Gu et al., 2024; Li et al., 2024; Schroeder & Wood-Doughty, 2024)
@@
After regeneration, we apply the reconstruction function again to obtain the updated semantic anchors $\widehat{S}_{\mathrm{sem}}^{(t+1)}$, and repeat the consistency check. This loop continues until the error set is empty or the maximum number of iterations $T_{\max}$ is reached. A more detailed discussion of the convergence is provided in the Appendix A.4. Upon termination, we return the final executable model $\mathcal{M}$, which combines the deterministic components with the latest semantically aligned code. The complete procedure of SAC-Opt is summarized in Algorithm 1.
@@ 4.4 Overall Performance
Reflexion 68.2% 49.2% 85.8% 43.2% 82.4% 76.1% 42.6%
OptiMUS-0.2 69.2% 44.0% 89.2% 45.8% 86.5% 75.8% 48.9%
OptiMUS-0.3 79.8% 54.0% 92.4% 52.1% 89.8% 81.0% 52.2%
SAC-Opt 86.8% 63.7% 96.5% 79.6% 94.0% 88.7% 58.9%
Impr. 7.0% $\uparrow$ 9.7% $\uparrow$ 2.1% $\uparrow$ 21.9% $\uparrow$ 4.2% $\uparrow$ 7.7% $\uparrow$ 1.8% $\uparrow$
@@ 4.6 Generalization Evaluation
Although the performance naturally depends on the reasoning ability of the underlying LLM, we emphasize that SAC-Opt is model-agnostic by design. Its modular pipeline, consisting of semantic anchor extraction, semantic verification, and correction, can be instantiated with any sufficiently capable model. To assess generalization of the proposed correction method, we further evaluated SAC-Opt on the open-source Qwen2.5-72B-Instruct model while keeping the extraction step fixed for fairness. As shown in Table 3, although Qwen2.5-72B-Instruct achieves lower base accuracy than GPT-4o, SAC-Opt consistently delivered clear gains across all datasets. This shows that even with a less powerful model, our correction mechanism still provides substantial benefits. These findings confirm that the effectiveness of SAC-Opt stems from its semantic correction mechanism rather than reliance on any specific model, demonstrating robust generalization across different LLMs.
@@ 5 Conclusion
We presented SAC-Opt, a backward semantic-anchored correction framework for optimization modeling that explicitly addresses semantic inconsistencies in LLM-based models. By aligning reconstructed anchors from generated models with the original task description, SAC-Opt iteratively corrects only the mismatched components, driving convergence toward semantically faithful solutions. This backward, anchor-guided refinement moves beyond solver-driven syntactic checks, enabling fine-grained correction of constraints and objective without additional training or supervision. Experiments on seven public datasets demonstrate an average accuracy gain of 7.8%. These findings underscore the reliability of our semantic-anchored correction framework for LLM-based optimization workflows.
@@ A.4 Discussion of Convergence
Accuracy Best-baseline 79.8% 54.0% 92.4% 52.1% 89.8% 81.0% 52.2%
SAC-Opt 86.8% 63.7% 96.5% 79.6% 94.0% 88.7% 58.9%
Impr. 7.0 $\uparrow$ 9.7 $\uparrow$ 2.1 $\uparrow$ 21.9 $\uparrow$ 4.2 $\uparrow$ 7.7 $\uparrow$ 1.8 $\uparrow$
@@
SAC-Opt 78.43 79.00 92.88 40.96 73.97 79.85 42.02
Diff. 13.76 $\uparrow$ 34.08 $\downarrow$ 4.09 $\uparrow$ 32.74 $\uparrow$ 2.11 $\uparrow$ 5.62 $\uparrow$ 26.66 $\downarrow$

### v2

# SAC-Opt: Semantic Anchors for Iterative Correction in Optimization Modeling

@@ Abstract
Large language models (LLMs) have opened new paradigms in optimization modeling by enabling the generation of executable solver code from natural language descriptions. Despite this promise, existing approaches typically remain solver-driven: they rely on single-pass forward generation and apply limited post-hoc fixes based on solver error messages, leaving undetected semantic errors that silently produce syntactically correct but logically flawed models. To address this challenge, we propose SAC-Opt, a backward-guided correction framework that grounds optimization modeling in problem semantics rather than solver feedback. At each step, SAC-Opt aligns the original semantic anchors with those reconstructed from the generated code and selectively corrects only the mismatched components, driving convergence toward a semantically faithful model. This anchor-driven correction enables fine-grained refinement of constraint and objective logic, enhancing both fidelity and robustness without requiring additional training or supervision. Empirical results on seven public datasets demonstrate that SAC-Opt improves average modeling accuracy by 7.7%, with gains of up to 21.9% on the ComplexLP dataset. These findings highlight the importance of semantic-anchored correction in LLM-based optimization workflows to ensure faithful translation from problem intent to solver-executable code.
@@ 1 Introduction
Optimization problems arise across domains such as logistics, healthcare, and finance, supporting tasks like planning, allocation, and portfolio optimization (Antoniou and Lu, 2007; Singh, 2012). These problems are typically formulated as mathematical programs and solved using external solvers such as Gurobi (Bixby, 2007), CPLEX (Cplex, 2009), or COPT Ge et al. (2023). However, translating real-world scenarios into solver-executable code often requires collaboration between domain experts and engineers. This process is time-consuming, hard to scale, and largely inaccessible to non-experts, as reflected by a survey showing that 81% of Gurobi users hold advanced degrees, with nearly half specializing in operations research (Optimization, 2023).
@@
(1) We introduce SAC-Opt, the first optimization modeling framework that performs proactive semantic verification to detect silent semantic errors that solver-driven checks cannot capture. (2) We propose a backward, semantic anchor-guided correction mechanism that progressively aligns models with problem intent, achieving convergence through fine-grained refinement. (3) We evaluate SAC-Opt on seven public datasets and show that it improves modeling accuracy by 7.7% on average, with a 21.9% gain on the challenging ComplexLP dataset.
@@ 3.5 Iterative Semantic Alignment and Correction
LLM-based Verification: (Gu et al., 2024; Li et al., 2024; Schroeder and Wood-Doughty, 2024)
@@
After regeneration, we apply the reconstruction function again to obtain the updated semantic anchors $\widehat{S}_{\mathrm{sem}}^{(t+1)}$, and repeat the consistency check. This loop continues until the error set is empty or the maximum number of iterations $T_{\max}$ is reached. A more detailed discussion of the convergence is provided in the Appendix A.4. Upon termination, we return the final executable model $\mathcal{M}$, which combines the deterministic components $\mathcal{M}_{\mathrm{simp}}$ with the latest semantically aligned code $\mathcal{M}_{\mathrm{sem}}^{(t)}$. The complete procedure of SAC-Opt is summarized in Algorithm 1.
@@ 4.4 Overall Performance
Reflexion 68.2% 49.0% 85.8% 43.2% 82.4% 76.1% 42.2%
OptiMUS-0.2 69.2% 43.8% 89.2% 45.8% 86.5% 75.8% 48.9%
OptiMUS-0.3 79.8% 54.3% 92.4% 52.1% 89.8% 81.0% 52.2%
SAC-Opt 86.8% 63.8% 96.5% 79.6% 94.0% 88.7% 58.9%
Impr. 7.0% $\uparrow$ 9.5% $\uparrow$ 2.1% $\uparrow$ 21.9% $\uparrow$ 4.2% $\uparrow$ 7.7% $\uparrow$ 1.8% $\uparrow$
@@ 4.6 Generalization Evaluation
Although the performance naturally depends on the reasoning ability of the underlying LLM, our goal here is to examine whether the benefits of SAC-Opt come from the semantic correction mechanism itself rather than from the strength of any particular model. The pipeline of SAC-Opt, consisting of semantic anchor extraction, semantic verification, and correction, can be instantiated with different LLMs as long as they have adequate reasoning capacity. To assess this, we further evaluated SAC-Opt using the open-source Qwen2.5-72B-Instruct model while keeping the extraction step fixed for fairness. As shown in Table 3, although Qwen2.5-72B-Instruct yields lower base accuracy than GPT-4o, SAC-Opt consistently provides clear improvements across all datasets. This indicates that the gains stem from the semantic correction process rather than dependence on a specific LLM, and that SAC-Opt remains effective even when paired with a less capable model.
@@ 5 Conclusion
We presented SAC-Opt, a backward semantic-anchored correction framework for optimization modeling that explicitly addresses semantic inconsistencies in LLM-based models. By aligning reconstructed anchors from generated models with the original task description, SAC-Opt iteratively corrects only the mismatched components, driving convergence toward semantically faithful solutions. This backward, anchor-guided refinement moves beyond solver-driven syntactic checks, enabling fine-grained correction of constraints and objective without additional training or supervision. Experiments on seven public datasets demonstrate an average accuracy gain of 7.7%. These findings underscore the reliability of our semantic-anchored correction framework for LLM-based optimization workflows.
@@ A.4 Discussion of Convergence
Accuracy Best-baseline 79.8% 54.3% 92.4% 52.1% 89.8% 81.0% 57.1%
SAC-Opt 86.8% 63.8% 96.5% 79.6% 94.0% 88.7% 58.9%
Impr. 7.0 $\uparrow$ 9.5 $\uparrow$ 2.1 $\uparrow$ 21.9 $\uparrow$ 4.2 $\uparrow$ 7.7 $\uparrow$ 1.8 $\uparrow$
@@
SAC-Opt 78.43 80.20 92.88 40.96 73.97 79.85 42.02
Diff. 13.76 $\uparrow$ 32.88 $\downarrow$ 4.09 $\uparrow$ 32.74 $\uparrow$ 2.11 $\uparrow$ 5.62 $\uparrow$ 26.66 $\downarrow$